# IFT UAM-CSIC 24-183 DESY-24-205

### v2

@@ Appendix A The bispectrum
where $-\infty_{-}=-\infty(1-i\omega)$ is responsible for projecting onto the interaction vacuum and the coefficients $C_{n}$ can be expressed in terms of the coefficients $c_{n}$, defined in Eq. (29), via
$C_{0}(t)$ $\equiv 6c_{0}(t)a(t)^{3}H(t)^{2}\varphi_{p}(t)\varphi_{q}(t)\varphi_{k}(t),$ (66)
$C_{1}(t)$ $\equiv 2c_{1}(t)a(t)^{3}\varphi_{p}(t)\dot{\varphi}_{q}(t)\dot{\varphi}_{k}(t)+(p\leftrightarrow q,k),$ (67)
$C_{2}(t)$ $\equiv c_{2}(t)a(t)^{3}\dot{\varphi}_{p}(t)\bigg{[}\frac{{\bm{q}}\cdot{\bm{k}}}{k^{2}}\varphi_{q}(t)\dot{\varphi}_{k}(t)+(q\leftrightarrow k)\bigg{]}+(p\leftrightarrow q,k),$ (68)
$C_{3}(t)$ $\equiv-2c_{3}(t)a(t)({\bm{q}}\cdot{\bm{k}})\varphi_{p}(t)\varphi_{q}(t)\varphi_{k}(t)+(p\leftrightarrow q,k),$ (69)
$C_{4}(t)$ $\equiv 2c_{4}(t)a(t)^{3}\frac{({\bm{q}}\cdot{\bm{k}})^{2}}{q^{2}k^{2}}\varphi_{p}(t)\dot{\varphi}_{q}(t)\dot{\varphi}_{k}(t)+(p\leftrightarrow q,k),$ (70)
@@
$A\equiv 6a^{3}H^{2}c_{0}+2p^{2}a\,c_{3},\qquad B\equiv 2a^{3}c_{1}+2a^{3}c_{4},\qquad C\equiv 4a^{3}c_{1}-2a^{3}c_{2}+4\frac{({\bm{k}}\cdot{\bm{p}})^{2}}{k^{2}p^{2}}a^{3}c_{4}.$ (72)
@@
$\frac{\mathcal{I}(k,p)}{|\varphi_{k}(t)|^{2}|\varphi_{p}(t)|^{2}}=-\epsilon(t)-2\eta(t)+2\,{\rm Im}\bigg{\{}\frac{\varphi_{p}(t)^{2}}{|\varphi_{p}(t)|^{2}}\int_{-\infty_{-}}^{t}a^{3}H^{2}\bigg{[}\mathcal{C}(t^{\prime})-4\epsilon^{2}\bigg{(}\frac{p}{aH}\bigg{)}^{2}\bigg{]}\varphi_{p}^{*2}(t^{\prime})\mathop{}\!\mathrm{d}t^{\prime}\bigg{\}},$ (74)
@@
${\rm Im}\bigg{[}\frac{\varphi_{p}(t)^{2}\varphi_{p}^{*}(t^{\prime})^{2}}{|\varphi_{p}(t)|^{2}}\bigg{]}=-\dfrac{1}{2}\int_{t^{\prime}}^{t}\left(\dfrac{1}{a^{3}\epsilon}\right){\rm d}t^{\prime\prime}+\mathcal{O}\bigg{[}\dfrac{p}{a(t)H(t)}\bigg{]}\,.$ (77)
@@ Appendix B Failure of the Gram-Charlier series
As a simple example of a situation in which the above series does not converge, we can use the PDF in Eq. (23). We choose $\sigma=0.2$ and $\eta_{\rm CR}=-2$. We find the coefficients via Eq. (83) and show the resulting PDF truncated order by order in Figure 7. It is apparent that the convergence of the series becomes progressively worse as more terms are added. Indeed, the Gram-Charlier series is, in general, not an appropriate way to reconstruct a PDF unless the latter is close enough to a Gaussian.

### v3

@@ Appendix A The bispectrum
where $\infty_{-}=-\infty(1-i\omega)$ is responsible for projecting onto the interaction vacuum and the coefficients $C_{n}$ can be expressed in terms of the coefficients $c_{n}$, defined in Eq. (29), via
$C_{0}(t)$ $\equiv 6M_{P}^{2}\,c_{0}(t)a(t)^{3}H(t)^{2}\varphi_{p}(t)\varphi_{q}(t)\varphi_{k}(t),$ (66)
$C_{1}(t)$ $\equiv 2M_{P}^{2}\,c_{1}(t)a(t)^{3}\varphi_{p}(t)\dot{\varphi}_{q}(t)\dot{\varphi}_{k}(t)+(p\leftrightarrow q,k),$ (67)
$C_{2}(t)$ $\equiv M_{P}^{2}\,c_{2}(t)a(t)^{3}\dot{\varphi}_{p}(t)\bigg[\frac{{\bm{q}}\cdot{\bm{k}}}{k^{2}}\varphi_{q}(t)\dot{\varphi}_{k}(t)+(q\leftrightarrow k)\bigg]+(p\leftrightarrow q,k),$ (68)
$C_{3}(t)$ $\equiv-2M_{P}^{2}\,c_{3}(t)a(t)({\bm{q}}\cdot{\bm{k}})\varphi_{p}(t)\varphi_{q}(t)\varphi_{k}(t)+(p\leftrightarrow q,k),$ (69)
$C_{4}(t)$ $\equiv 2M_{P}^{2}\,c_{4}(t)a(t)^{3}\frac{({\bm{q}}\cdot{\bm{k}})^{2}}{q^{2}k^{2}}\varphi_{p}(t)\dot{\varphi}_{q}(t)\dot{\varphi}_{k}(t)+(p\leftrightarrow q,k),$ (70)
@@
$\frac{A}{M_{P}^{2}}\equiv 6a^{3}H^{2}c_{0}+2p^{2}a\,c_{3},\qquad\frac{B}{M_{P}^{2}}\equiv 2a^{3}c_{1}+2a^{3}c_{4},\qquad\frac{C}{M_{P}^{2}}\equiv 4a^{3}c_{1}-2a^{3}c_{2}+4\frac{({\bm{k}}\cdot{\bm{p}})^{2}}{k^{2}p^{2}}a^{3}c_{4}.$ (72)
@@
$\frac{\mathcal{I}(k,p)}{|\varphi_{k}(t)|^{2}|\varphi_{p}(t)|^{2}}=-\epsilon(t)-2\eta(t)+2M_{P}^{2}\,{\rm Im}\bigg\{\frac{\varphi_{p}(t)^{2}}{|\varphi_{p}(t)|^{2}}\int_{-\infty_{-}}^{t}a^{3}H^{2}\bigg[\mathcal{C}(t^{\prime})-4\epsilon^{2}\bigg(\frac{p}{aH}\bigg)^{2}\bigg]\varphi_{p}^{*2}(t^{\prime})\mathop{}\!\mathrm{d}t^{\prime}\bigg\},$ (74)
@@
${\rm Im}\bigg[\frac{\varphi_{p}(t)^{2}\varphi_{p}^{*}(t^{\prime})^{2}}{|\varphi_{p}(t)|^{2}}\bigg]=-\dfrac{1}{2M_{P}^{2}}\int_{t^{\prime}}^{t}\left(\dfrac{1}{a^{3}\epsilon}\right){\rm d}t^{\prime\prime}+\mathcal{O}\bigg[\dfrac{p}{a(t)H(t)}\bigg]\,.$ (77)
@@ Appendix B Failure of the Gram-Charlier series
As a simple example of a situation in which the above series does not converge, we can use the PDF in Eq. (23). We choose $\sigma=0.2$ and $\eta_{\rm CR}=-2$. We find the coefficients via Eq. (83) and show the resulting PDF truncated order by order in Figure 7. It is apparent that the convergence of the series becomes progressively worse as more terms are added. Indeed, the Gram-Charlier series is, in general, not an appropriate way to reconstruct a PDF unless the latter is close enough to a Gaussian. Notice that the PDF from the Gram-Charlier series even becomes locally negative at certain perturbative orders, something that by construction does not happen with the procedure we have put forward in this work.